OpenStudy (anonymous):

Find the length of the hypotenuse. A. 5 B. 6 C. 12 D. 18

OpenStudy (anonymous):



OpenStudy (caozeyuan):

hypo=sqrt(2)*leg! REMEMBER THAT!

OpenStudy (caozeyuan):

leg=sqrt(2)*3 so hypo=3*2=6

OpenStudy (anonymous):

thank you.